# Eccentricity signatures in LIGO-Virgo-KAGRA’s BNS and NSBH binaries

### v1

@@ Abstract
[…]
Gravitational Waves, Compact Binaries
@@ I Introduction
Advanced LIGO and Advanced Virgo [1, 2] have detected hundreds of gravitational wave (GW) signals from mergers of compact objects during their third observing run (O3), covering a wide range of source masses [3, 4, 5, 6, 7]. Although the majority of these events are binary black hole (BBH) mergers, a handful of low-mass events fall within the binary neutron star (BNS) and neutron star–black hole (NSBH) mass ranges. These include GW170817 [8], GW190425 [9] identified as BNS mergers, and GW200105, GW200115 [10], and the recently observed GW230529 [11] in the fourth observing run of LIGO and Virgo, as NSBH mergers based primarily on their component masses. These detections have enabled the observations and constraints of several known features of compact binary systems, such as spin-induced precession of the binary’s orbit in GW190412 [12] and the presence of subdominant harmonics in GW190412 and GW190814 [12, 13].
@@
Until the end of O3, there was no compelling observational evidence for orbital eccentricity in compact binary mergers in the low-mass systems such as NSBHs, BNSs, or even the lowest mass BBHs. However, recent analyses [14, 15, 16] of GW200105 have indicated possible hints of eccentricity, marking it as the first NSBH event with potential eccentricity. This observation challenges the assumption of an isolated binary origin and suggests a formation channel beyond standard binary evolution models [17, 18, 19].
@@
Eccentricity leaves a characteristic imprint on the gravitational-wave signal by modulating the phase evolution during the early inspiral. As binary orbits, the orbital eccentricity dissipates rapidly and generally circularizes as the system approaches the merger phase unless acted upon by external forces [20]. However, if residual eccentricity persists in the frequency band of ground-based detectors like LIGO and Virgo (around 102010-2010 - 20 Hz), it can become detectable. Since binary’s orbital eccentricity carries traces of its evolutionary history, measuring it in a GW signal can serve as a smoking gun indicator for its distinct formation channel [18].
@@
A compact binary system can exhibit measurable orbital eccentricity if it is either born with high eccentricity or acquired through external interactions over time. The first scenario is more common for isolated stellar-mass binary stars that evolve through pathways such as the common envelope phase [21, 22, 23, 24], chemically homogeneous evolution [25, 26], or fallback-driven mechanisms in failed supernovae [27]. In these cases, eccentricity is unlikely to be detectable, as it is expected to dampen to $e\lesssim 10^{-3}$ when the GW signal emitted enters the LIGO-Virgo frequency band ($>10\,\rm Hz$[28, 29]. The second scenario typically occurs in dense environments, such as nuclear star clusters, young star clusters, and globular clusters, in which BHs form from the collapse of massive stars [30, 31]. In these regions, dynamical interactions, such as frequent gravitational encounters and dynamic friction [32, 33], can induce measurable eccentricity in the orbit of compact binaries.
@@
Although dynamical formation channels predict large eccentricities for high-mass binaries and have been extensively studied in the literature [34, 35, 36, 37, 38, 39, 40], there has been growing interest in exploring eccentricity in low-mass systems as well due to the recent support for eccentricity in GW200105 [14, 15, 16]. However, inferring properties from these systems is challenging due several computational challenges. First, including eccentricity into the search space significantly increases dimensionality and parameter ranges, making matched filtering searches much more computationally expensive [41]. Second, eccentricity is most pronounced during the early inspiral stages, where the LIGO and Virgo detector’s poor low-frequency sensitivity ($<10\,\rm Hz$) limits their detectability. Third, searching for these features at lower frequencies extends the signal duration, further increasing the computational cost of Bayesian inference. Finally, the challenge is compounded by the lack of accurate waveform models, making such searches or Bayesian inferences even more difficult.
@@
In this study, we present a systematic investigation of orbital eccentricity in low-mass gravitational-wave events observed by LIGO and Virgo up to the publicly available portion of the fourth observing run. Our analysis focuses on inferring eccentricity using the SEOBNRv5EHM model, an effective-one-body (EOB) formalism [42, 43] built over SEOBNRv5 waveform family [44, 45, 46, 47, 48] that incorporates higher-order modes and eccentricity effects [49, 50]. For comparison, we also use the TEOBResumS-Dali model [51] as an alterative EOB-based waveform model to assess the consistency and accuracy of eccentricity constraints across models. Our analysis is performed over the dominant (2,2) mode of the waveform, and as well as the higher order modes (HOM) up to $l=4$, which include subdominant multi-polar contributions beyond the quadrupole that become important in systems such as high mass ratios or eccentric orbits.
@@
We find that five out of the six events are consistent with circular orbits, placing stringent upper bounds on eccentricity. However, for GW200105, we observe moderate support for non-zero eccentricity across both waveform models. In particular, GW200105 exhibits a well-constrained eccentricity of $0.125^{+0.029}_{-0.082}$ for SEOBNRv5EHM and $0.135^{+0.019}_{-0.088}$ for TEOBResumS-Dali with 90% credible level with the inclusion of HOMs and a reference frequency of 20 Hz, consistent with previous studies [14, 15, 16].
@@
To better understand the underlying structure of the posterior distributions and sampling challenges, we also conduct a thorough investigation of the eccentricity–anomaly parameter space. Our analysis reveals multi-modal features and isolated regions in the posterior distributions, particularly for systems with non-negligible eccentricity such as GW200105. We perform multiple checks, including varying the frequency cutoff across multiple inferences, injecting signals at higher frequency values, calculating the likelihoods within the peaks and dips of the modes, and using alternate samplers. Furthermore, by directly computing waveform overlaps while varying only the eccentricity, we demonstrate that these modes arise from the waveform model itself and not from the sampling procedure. This behavior is consistent across both TEOBResumS-Dali and SEOBNRv5EHM and suggests that the observed structure is physical and not due to waveform systematics.
@@ II Method
The posterior distribution represents the probability of parameters $\vec{\theta}$ given the observed data $d$ and the model $h$ [55, 56]. The likelihood $\mathcal{L}(d|\vec{\theta},h)$ quantifies how well the data match a particular set of parameters and the waveform model. $\pi(\vec{\theta}|h)$ is the prior distribution for a set of parameters. The prior represents our initial assumptions or beliefs about the ranges of the parameters, based on existing knowledge or theoretical considerations. $\mathcal{Z}$ is the evidence and serves as a normalization factor in this case. $\vec{\theta}$ represents the intrinsic and extrinsic parameters of the binary system. Intrinsic parameters can include the masses of the compact objects, their individual spins, the tidal deformability of the compact object if it is a neutron star, the orbital eccentricity and the radial phase parameter, which characterizes the binary’s relative position within a Keplerian orbit. The radial phase can be expressed either as the mean anomaly or the relativistic anomaly. The extrinsic parameters consist of the distance from the source, its location, the angle of inclination to the detector, the angle of polarization with respect to the detector, and the time and phase of the binary coalescence.
@@ II.1 Waveform Model
To measure eccentricity in low-mass gravitational-wave events, we employ two time-domain waveform models within the EOB framework SEOBNRv5EHM [57, 49] and TEOBResumS-Dali [58]. Both models generate HOM eccentric inspiral–merger–ringdown waveforms and are calibrated with numerical relativity simulations. The SEOBNRv5EHM model is built for binaries with eccentric orbits and spins aligned (or antialigned) with the orbital angular momentum. In addition to the dominant (2,2) mode, it includes subdominant modes such as (2,1), (3,3), (3,2), (4,4), and (4,3), allowing for more accurate modeling of systems with mass asymmetry or orbital inclination. It is important to note that this model utilizes the definition of relativistic anomaly to describe the position of the binary in the eccentric orbit.
TEOBResumS-Dali is another time-domain waveform model with the ability to model a wide range of compact binary systems, including BNS, BBH, and NSBH binaries [58]. It incorporates a comprehensive set of physical effects critical for accurate waveform modeling, including orbital eccentricity, hyperbolic encounters, non-planar orbital dynamics, spin-orbit and spin-spin couplings, and the tidal deformability of neutron stars. Additionally, the model includes HOMs up to the multipole order $l=8$. Instead of relativistic anomaly, TEOBResumS-Dali uses the mean anomaly definition to describe the radial phase of the eccentric orbit.
Since both models are not originally formulated in the frequency domain, we apply a Fast Fourier Transform with appropriate tapering to convert the waveform for use in our frequency-domain inference pipeline. Before performing the inference, we validate that the waveform was generated successfully across our parameter space and that all tested waveforms were produced without issue.
@@ II.2 Data and Priors
We analyze six events—GW170817, GW190425, GW190814, GW200105, GW200115, and GW230529, using publicly available strain data from the LIGO Hanford, LIGO Livingston, and Virgo detectors [59]. For each event, we include data only from the detectors that were online at the time of the observation. GW190425 and GW200105 were recorded with the LV network, and GW230529 was observed only by LIGO Livingston. Although both Livingston and Virgo data were available for GW200105, we excluded Virgo from our analysis. Virgo’s lower sensitivity and unfavorable sky location significantly reduce it’s contribution to the signal, particularly in the lower frequency band relevant for an eccentric study. For computational efficiency and minimizing excessive artifacts, we only use Livingston data in our final analysis for GW200105.
We choose prior bounds that align with the astrophysical properties measured in the Fourth Open Gravitational Wave Catalog [4] and the Fourth Observation Run Discovery Paper with the exception of precessing spins [60]. Specifically, we choose uniform priors over the detector frame chirp mass $\mathcal{M}$, the mass ratio $q$, and eccentricity $e$. We choose the bounds on the eccentricity to range from 0 to 0.2 for GW200105 and 0 to 0.1 for the rest of the sources. Additionally, we perform a separate inference using a uniform-in-log prior on eccentricity for GW200105 to assess the dependence of our results on the prior choice. For this run, we choose a minimum bound of $10^{-4}$ and an upper bound of 0.2 Incorporating eccentricity, HOM, and precession into gravitational waveform models remains an area of ongoing development, and current models have varying levels of support for these effects. Due to existing model limitations, we adopt a broad agnostic prior with uniform aligned spin, $\chi_{1z}$ and $\chi_{2z}$. We choose uniform angle priors for the polarization $\Psi$, relativistic anomaly $\ell_{\text{rel}}$ for SEOBNRv5EHM and mean anomaly $\ell_{m}$ for TEOBResumS-Dali. We choose a uniform luminosity volume prior for $d_{L}$. We assumed an isotropic sky distribution for the inclination $\iota$, right ascension $\alpha$, and declination $\delta$. In our study, we do not impose any prior on the tidal deformability parameter of neutron stars. The reason is that we do not expect the tidal effects to be measurable with the current detector’s sensitivities for the given events. As a result, its inclusion is unlikely to affect the measurement of the eccentricity parameter.
@@ II.3 Parameter Estimation Optimizations
Bayesian inference for low-mass binaries is computationally demanding, primarily due to the long-duration signals and the slow waveform generation associated with EOB models. For example, generating a waveform from our highest likelihood model for GW190425 took 12.3 seconds to generate in the frequency domain. During parameter estimation, $\sim\mathcal{O}(10^{6})$ likelihoods are calculated, making parameter estimation for these systems especially costly. Although reduced order models are often used in analyses to speed up EOB waveform evaluation, no such surrogate currently exists for the parameter space and models we utilized [61, 62, 63]. As a result, we directly use the complete EOB models, which require the numerical solution of coupled ordinary differential equations to compute the waveform dynamics [64, 58]. This waveform generation cost is further amplified when modeling complex systems such as eccentric orbits with HOMs. In addition, the high dimensionality and complex structure of the parameter space substantially increase the computational cost of efficient sampling and can lead to slow convergence or inefficient exploration of the posterior.
@@
Given that the computational cost of Bayesian inference is expected to be large for these signals, we employ a few optimizations. One optimization includes reducing the dimensionality of our parameter space to make sampling more efficient. To do this, we employ marginalization techniques [54, 65] where we integrate the time of coalescence, sky location, distance, and polarization angle from the likelihood calculation. This approach improves sampling efficiency and convergence, thereby accelerating the inference process. Marginalization over distance is performed using a precomputed lookup table, while sky location, time of arrival, and polarization are sampled from a proposal distribution approximating the posterior [65, 66]. The samples are then reweighted by the ratio of the posterior, following the method outlined in Ref. [65]. Marginalization over the coalescence phase is also feasible. However, this method can only be done for the dominant (2,2) mode waveforms, as marginalization techniques over the coalescence phase are not currently implemented for waveforms with HOMs and remains an ongoing area of development. As a result, the phase must be explicitly sampled, increasing the dimensionality of the parameter space and further adding to the computational costs.
@@
For events with non-negligible eccentricity, we also utilize a parallelization strategy by splitting the prior range of eccentricity evenly into multiple bins, running independent inference in each and recombining in post. This approach helps mitigate the complexity of the likelihood surface: eccentric binaries introduce strong, nonlinear structure in the posterior, which can be highly multi-modal and slow to converge when explored globally. By dividing the prior into narrower intervals, each sampler explores a smaller, more manageable region of parameter space, improving convergence and reducing overall computational cost.
@@
To efficiently explore our high-dimensional parameter space, we utilize the nested sampler Dynesty [67]. Nested sampling is particularly well-suited for handling complex, multi-modal distributions and efficiently sampling high-dimensional likelihood surfaces. This is especially beneficial in our analysis, where the incorporation of both eccentricity and HOM significantly increases the complexity of the likelihood function.

### v2

# Eccentricity signatures in LIGO-Virgo-KAGRA’s Binary Neutron stars and Neutron-star Black Holes

@@ Abstract
Measurement of the eccentricity of low-mass binary systems through gravitational waves is crucial to distinguish between various formation channels. However, detecting eccentricity in these systems has been challenging due to the lack of accurate eccentric waveform models and the high computational cost associated with Bayesian inferences for systems with low-mass objects. In this work, we assess the eccentricities of seven previously observed low-mass gravitational wave events using publicly available data from the O1-O4a observing runs of the LIGO and Virgo observatories. We analyze the events using a new eccentric waveform model, SEOBNRv5EHM, and compare our results with the existing model TEOBResumS-Dali. We also present the first eccentricity constraints for GW190814. To improve the accuracy of our parameter estimation, we include higher-order modes in both waveform models. We optimize inference by employing efficient marginalization techniques to alleviate the computational costs associated with low-mass systems, and parallelization techniques for sampling large parameter spaces. We find that one of the sources, GW200105, exhibits a non-negligible eccentricity, with a measured eccentricity of $e=0.125^{+0.029}_{-0.082}$ at 20 Hz ($90\%$ credible level) for SEOBNRv5EHM and $e=0.135^{+0.019}_{-0.088}$ for TEOBResumS-Dali for a uniform eccentricity prior ranging from 0 to 0.2 at a reference frequency of 20 Hz. We find moderate support for the eccentric waveform hypothesis with a Bayes factor of $\sim 6-7$ times more preference for the eccentric model over the non-eccentric one. When using a uniform log prior on eccentricity with a minimum bound of $10^{-4}$, the support for the eccentric model decreases, with the Bayes factor reduced to $2.35$. For the remaining five sources, the results are consistent with low eccentricity, with 90% upper limits ranging from $e\leq 0.011$ to $e\leq 0.066$. We do not find any support for non-negligible eccentricity in GW190814. Finally, we discuss the challenges of performing Bayesian inference in eccentric, multi-modal parameter spaces, including issues related to sampling efficiency and waveform systematics.
Gravitational Waves, Compact Binaries, Eccentric Binaries
@@ I Introduction
Advanced LIGO and Advanced Virgo [1, 2] have detected hundreds of gravitational wave (GW) signals from the mergers of compact objects during their third observing run (O3), covering a wide range of source masses [3, 4, 5, 6, 7]. Although the majority of these events are binary black hole (BBH) mergers, a handful of low-mass events fall within the binary neutron star (BNS) and neutron star–black hole (NSBH) mass ranges. These include GW170817 [8], GW190425 [9] identified as BNS mergers, and GW200105, GW200115 [10], and the two recently observed objects GW230529 and GW230518 [11, 12, 13] in the first part of the fourth observing run (O4a) of LIGO, identified as NSBH mergers based primarily on their component masses. These detections have enabled observations and constraints on several known features of compact binary systems, such as the spin-induced precession of the binary’s orbit in GW190412 [14, 15, 16] and the presence of subdominant harmonics in GW190412 and GW190814 [14, 17].
@@
Until the end of O3, there was no compelling observational evidence for orbital eccentricity in compact binary mergers in low-mass systems such as NSBHs, BNSs, or even low-mass BBHs. However, recent analyses of GW200105 have indicated possible hints of eccentricity, marking it as the first NSBH event with potential eccentricity [18, 19, 20]. This observation challenges the assumption of an isolated binary origin and suggests a formation channel beyond standard binary evolution models [21, 22, 23].
@@
Eccentricity leaves a characteristic imprint on the gravitational-wave signal by modulating the phase evolution during the early inspiral. As the binary orbits, the orbital eccentricity dissipates rapidly and generally circularizes as the system approaches the merger phase unless acted upon by external forces [24]. However, if residual eccentricity persists in the frequency band of ground-based detectors like LIGO and Virgo (around 102010-20 Hz), it can be detectable. Since the binary’s orbital eccentricity carries traces of its evolutionary history, measuring it in a GW signal can serve as a smoking gun indicator for its distinct formation channel [22].
@@
A compact binary system can exhibit measurable orbital eccentricity if it is either born with high eccentricity or acquires it through external interactions over time. The first scenario is more common for isolated stellar-mass binary stars that evolve through pathways such as the common envelope phase [25, 26, 27, 28], chemically homogeneous evolution [29, 30], or fallback-driven mechanisms in failed supernovae [31]. In these cases, eccentricity is unlikely to be detectable, as it is expected to dampen to $e\lesssim 10^{-3}$ when the GW signal emitted enters the LIGO-Virgo frequency band ($10\,\rm Hz$) [32, 33]. The second scenario typically occurs in dense environments, such as nuclear star clusters, young star clusters, and globular clusters, in which binaries form and harden through dynamical interactions [34, 35]. In these regions, dynamical interactions, such as frequent gravitational encounters and dynamic friction [36, 37], can induce measurable eccentricity in the orbits of compact binaries.
@@
Although dynamical formation channels predict large eccentricities for high-mass binaries,theoretical studies generally find that the overall fraction of binaries with measurable eccentricity is small, typically a few percent, and even lower for low-mass systems [38, 39, 40]. This makes the observation of an eccentric low-mass binary even more unlikely [41, 42, 43, 44, 45, 46, 47]. Nevertheless, there has been growing interest in exploring eccentricity in low-mass systems as well, due to the recent support for eccentricity in GW200105 [18, 19, 20]. However, inferring properties from these systems is challenging due to several computational challenges. First, including eccentricity into the search space significantly increases dimensionality and parameter ranges, making matched filtering searches much more computationally expensive [48, 49, 50]. Second, eccentricity is most pronounced during the early inspiral stages, where the LIGO and Virgo detector’s poor low-frequency sensitivity ($10\,\rm Hz$) limits their detectability. Third, searching for these features at lower frequencies extends the signal duration, further increasing the computational cost of Bayesian inference. Finally, the challenge is compounded by the lack of accurate waveform models, making such searches or Bayesian inferences even more difficult.
@@
In this study, we present a systematic investigation of orbital eccentricity in low-mass gravitational-wave events observed by LIGO and Virgo between O1 and O4a. Our analysis focuses on inferring eccentricity using the SEOBNRv5EHM model, an effective-one-body (EOB) formalism [51, 52] built over SEOBNRv5 waveform family [53, 54, 55, 56, 57] that incorporates higher-order modes and eccentricity effects [58, 59]. For comparison, we also use the TEOBResumS-Dali model [60] as an alternative EOB-based waveform model to assess the consistency and accuracy of eccentricity constraints across models. Our analysis is performed over the dominant (2,2) mode of the waveform, as well as the higher-order modes (HOM) up to $l=4$, which include subdominant multi-polar contributions beyond the quadrupole that become important in systems such as high mass ratios or eccentric orbits.
@@
We find that six out of the seven events are consistent with circular orbits, placing stringent upper bounds on eccentricity. However, for GW200105, we observe moderate support for non-zero eccentricity across both waveform models. In particular, GW200105 exhibits a well-constrained eccentricity of $0.125^{+0.029}_{-0.082}$ for SEOBNRv5EHM and $0.135^{+0.019}_{-0.088}$ for TEOBResumS-Dali at the 90% credible level, including HOMs and a reference frequency of 20 Hz, consistent with previous studies [18, 19, 20].
@@
To better understand the underlying structure of the posterior distributions and the sampling challenges, we also conduct a thorough investigation of the eccentricity–anomaly parameter space. Our analysis reveals multi-modal features and isolated regions in the posterior distributions, particularly for systems with non-negligible eccentricity such as GW200105. We perform multiple checks, including varying the frequency cutoff across multiple inferences, injecting signals at higher frequency values, calculating the likelihoods within the peaks and dips of the modes, and using alternate samplers. Furthermore, by directly computing waveform overlaps while varying only the eccentricity, we demonstrate that these modes arise from the waveform model itself and not from the sampling procedure. This behavior is consistent in both SEOBNRv5EHM and TEOBResumS-Dali and suggests that the observed structure is physical and not due to waveform systematics.
@@ II Method
The posterior distribution represents the probability of the parameters $\vec{\theta}$ given the observed data $d$ and the model $h$ [64, 65]. The likelihood $\mathcal{L}(d|\vec{\theta},h)$ quantifies how well the data match a particular set of parameters and the waveform model. $\pi(\vec{\theta}|h)$ is the prior distribution for a set of parameters. The prior represents our initial assumptions or beliefs about the ranges of the parameters, based on existing knowledge or theoretical considerations. $\mathcal{Z}$ is the evidence and serves as a normalization factor in this case. $\vec{\theta}$ represents the intrinsic and extrinsic parameters of the binary system. Intrinsic parameters include the masses of the compact objects, their spins, the tidal deformability of the compact object if it is a neutron star, the orbital eccentricity, and the radial phase parameter, which characterizes the binary’s relative position within a Keplerian orbit. The radial phase can be expressed as either the true anomaly or the relativistic anomaly. The extrinsic parameters consist of the distance from the source, its location, the angle of inclination to the detector, the angle of polarization with respect to the detector, and the time and phase of the binary coalescence.
@@ II.1 Waveform Model
To measure eccentricity in low-mass gravitational-wave events, we employ two time-domain waveform models within the EOB framework: SEOBNRv5EHM [66, 58] and TEOBResumS-Dali [67]. Both models generate HOM eccentric inspiral–merger–ringdown waveforms and are calibrated with numerical relativity simulations. The SEOBNRv5EHM model is built for binaries with eccentric orbits and spins aligned (or antialigned) with the orbital angular momentum. In addition to the dominant (2,2) mode, it includes subdominant modes such as (2,1), (3,3), (3,2), (4,4), and (4,3), allowing for more accurate modeling of systems with mass asymmetry or orbital inclination. It is important to note that this model utilizes the definition of relativistic anomaly to describe the position of the binary in an eccentric orbit.
TEOBResumS-Dali is another time-domain waveform model with the ability to model a wide range of compact binary systems, including BNS, BBH, and NSBH binaries [67]. It incorporates a comprehensive set of physical effects critical for accurate waveform modeling, including orbital eccentricity, hyperbolic encounters, non-planar orbital dynamics, spin-orbit and spin-spin couplings, and the tidal deformability of neutron stars. Additionally, the model includes HOMs up to multipole order $l=8$. Instead of the relativistic anomaly, TEOBResumS-Dali uses the true anomaly definition to describe the radial phase of the eccentric orbit.
Since neither model is originally formulated in the frequency domain, we apply a Fast Fourier Transform with appropriate tapering to convert the waveform for use in our frequency-domain inference pipeline. Before performing the inference, we validate that the waveform was generated successfully across our parameter space and that all tested waveforms were produced without issues.
@@ II.2 Data and Priors
We analyze seven events—GW170817, GW190425, GW190814, GW200105, GW200115, GW230518, and GW230529 using publicly available strain data from the LIGO Hanford, LIGO Livingston, and Virgo detectors [68, 69, 70]. For each event, we include data only from the detectors that were online at the time of the observation. GW190425 and GW200105 were recorded with the Livingston-Virgo network, GW235018 [12, 13] with the Hanford-Livingston network, and GW230529 was observed only by LIGO Livingston. Although both Livingston and Virgo data were available for GW200105, we excluded Virgo from our analysis. Virgo’s lower sensitivity and unfavorable sky location significantly reduce its contribution to the signal, particularly in the lower frequency band relevant for an eccentric study. For computational efficiency and to minimize excessive artifacts, we only use the Livingston data in our final analysis for GW200105.
We choose prior bounds that align with the astrophysical properties measured in the Fourth Open Gravitational Wave Catalog [4] and the Fourth Observation Run Discovery Paper, except for precessing spins [71]. Specifically, we choose uniform priors over the detector frame chirp mass $\mathcal{M}$, the mass ratio $q$, and eccentricity $e$. We choose the bounds on the eccentricity to range from 0 to 0.2 for GW200105 and from 0 to 0.1 for the other sources. Additionally, we perform a separate inference using a uniform-in-log prior on eccentricity for GW200105 to assess the dependence of our results on the prior choice. For this run, we choose a minimum bound of $10^{-4}$ and an upper bound of 0.2. The reference frequency is set to be equal to the gravitational wave frequency of 20 Hz. Incorporating eccentricity, HOMs, and precession into gravitational waveform models remains an area of ongoing development, and current models have varying levels of support for these effects. Due to existing model limitations, we adopt a broad agnostic prior with uniform aligned spins, $\chi_{1z}$ and $\chi_{2z}$. We choose uniform angle priors for the polarization $\Psi$, relativistic anomaly $\ell_{\text{rel}}$ for SEOBNRv5EHM and true anomaly $\ell_{t}$ for TEOBResumS-Dali. We choose a uniform luminosity volume prior for $d_{L}$. We assumed an isotropic sky distribution for the inclination $\iota$, right ascension $\alpha$, and declination $\delta$. In our study, we do not consider the tidal deformability of neutron stars. The reason is that we do not expect the tidal effects to be measurable with the current detector’s sensitivity for the given events. As a result, its inclusion is unlikely to affect the measurement of the eccentricity parameter.
@@ II.3 Parameter Estimation Optimizations
Bayesian inference for low-mass binaries is computationally demanding, primarily due to the long-duration signals and the slow waveform generation associated with EOB models. For example, generating a waveform from our highest likelihood model for GW190425 took 12.3 seconds to generate in the frequency domain. During parameter estimation, $\sim\mathcal{O}(10^{6})$ likelihoods are calculated, making parameter estimation for these systems especially costly. Although reduced order models are often used in analyses to speed up EOB waveform evaluation, no such surrogate currently exists for the parameter space and models we utilized [72, 73, 74]. As a result, we directly use the complete EOB models, which require the numerical solution of coupled ordinary differential equations to compute the waveform dynamics [75, 67]. This waveform generation cost is further amplified when modeling complex systems, such as eccentric orbits with HOMs. In addition, the high dimensionality and complex structure of the parameter space substantially increase the computational cost of efficient sampling and can lead to slow convergence or inefficient exploration of the posterior.
@@
Given that the computational cost of Bayesian inference is expected to be large for these signals, we employ a few optimizations. One optimization includes reducing the dimensionality of our parameter space to make sampling more efficient. To do this, we employ marginalization techniques [63, 76] in which we integrate the time of coalescence, sky location, distance, and polarization angle from the likelihood calculation. This approach improves sampling efficiency and convergence, thereby accelerating the inference process. Marginalization over distance is performed using a precomputed lookup table, while sky location, time of arrival, and polarization are sampled from a proposal distribution approximating the posterior [76, 77]. The samples are then reweighted by the ratio of the posterior to the proposal, following the method outlined in Ref. [76]. Marginalization over the coalescence phase is also feasible. However, this method can only be applied to the dominant (2,2) mode waveforms, as marginalization techniques over the coalescence phase are not currently implemented for waveforms with HOMs and remain an ongoing area of development. As a result, the phase must be explicitly sampled, increasing the dimensionality of the parameter space and further adding to the computational costs.
@@
For events with non-negligible eccentricity, we also utilize a parallelization strategy by splitting the prior range of eccentricity evenly into multiple bins, running independent inference in each, and recombining in postprocessing. This approach helps mitigate the complexity of the likelihood surface: eccentric binaries introduce strong, nonlinear structure in the posterior, which can be highly multi-modal and slow to converge when explored globally. By dividing the prior into narrower intervals, each sampler explores a smaller, more manageable region of parameter space, improving convergence and reducing the overall computational cost.
@@
To efficiently explore our high-dimensional parameter space, we utilize the nested sampler Dynesty [78]. Nested sampling is particularly well-suited for handling complex, multi-modal distributions and efficiently sampling high-dimensional likelihood surfaces. This is especially beneficial in our analysis, where the incorporation of both eccentricity and HOMs significantly increases the complexity of the likelihood function.